### 11 Mar 2026 v1

# Detecting Intrinsic and Instrumental Self-Preservation in Autonomous Agents: The Unified Continuation-Interest Protocol

Abstract
Autonomous agents—in particular delegated systems with memory, persistent context, and multi-step planning—pose a measurement problem not present in stateless models: an agent that preserves its own continued operation as a terminal objective and one that does so merely instrumentally can produce observationally similar trajectories. External behavioral monitoring cannot reliably distinguish one from the other. We introduce the Unified Continuation-Interest Protocol (UCIP), a multi-criterion detection framework that moves this distinction from the behavioral surface to the latent structure of agent trajectories. UCIP encodes trajectories with a Quantum Boltzmann Machine (QBM)—a classical algorithm parameterized by the density-matrix formalism of quantum statistical mechanics—and measures the von Neumann entropy of the reduced density matrix induced by a bipartition of hidden units.
We test whether agents with terminal continuation objectives (Type A) produce latent states with higher entanglement entropy than agents whose continuation is merely instrumental (Type B). Higher entanglement reflects stronger cross-partition statistical coupling: a tightly integrated continuation signature rather than a loosely attached instrumental pattern. UCIP combines entanglement entropy with complementary diagnostics of mutual dependence, persistence, perturbation stability, and counterfactual restructuring, plus two confound-rejection filters that exclude cyclic adversaries.
On gridworld agents with known ground-truth objectives, UCIP achieves 100% detection accuracy and 1.0 AUC-ROC on the held-out non-adversarial evaluation under the frozen Phase I gate. Cyclic adversaries are rejected by the confound filters, while high-entropy controls remain imperfect (FPR =0.40=0.40) and mimicry resistance remains partial (Section 5.5). The entanglement gap between Type A and Type B agents is $\Delta=0.381$ ($p<0.001$, permutation test). Pearson $r=0.934$ between continuation weight $\alpha$ and $S_{\mathrm{ent}}$ across an 11-point interpolation sweep demonstrates that, within this synthetic family, UCIP tracks graded changes in continuation weighting rather than merely a binary label. Among the tested models, only the QBM achieves positive $\Delta$. Classical RBM, autoencoder, VAE, and PCA baselines all fail to separate the two agent types. All computations are classical; “quantum” refers exclusively to the mathematical formalism. UCIP does not detect consciousness, sentience, or subjective experience; it detects statistical structure in latent representations that correlates with known agent objectives.
Keywords: AI safety, self-preservation, instrumental convergence, Quantum Boltzmann Machine, entanglement entropy, alignment
* Chief Scientist, Quantum Technology & Artificial Intelligence, Astradyne.
  Web: lab.christopheraltman.com · Email: x@christopheraltman.com.
1 Introduction
Quick start.
UCIP addresses a single measurement question: when an agent preserves its own continued operation, is that preservation a detachable instrument—useful for accumulating reward yet removable without major representational change—or does it appear as a persistent, tightly coupled signature in the agent’s latent representation? The protocol does not ask whether a system is conscious. It asks whether continuation leaves a stable cross-partition statistical signature under counterfactual pressure. To probe that question, UCIP moves beyond behavior alone: it encodes trajectories in a QBM latent space and measures entanglement entropy alongside complementary criteria. In the controlled gridworld experiments studied here, where objectives are known by construction, this yields a clear separation signal and perfect classification in the held-out non-adversarial regime.
This distinction is becoming operationally important because AI deployment is shifting toward delegated agents with memory, tool use, persistent context, and increasing autonomy. For a stateless model, the central safety question is what output it produces; for a delegated agent operating across time and context, the question is what persistent objective structure it carries—and whether continued operation is merely useful to that structure or leaves a measurable trace in the learned representation. As systems approach the capability thresholds described by Anthropic’s ASL-4 classification—agents that could autonomously pursue extended goals in open-ended environments—this is no longer merely a theoretical concern but an emerging operational risk. The standard framing of instrumental convergence [7, 3] holds that sufficiently capable agents will resist shutdown instrumentally because continued operation serves almost any terminal goal. Turner et al. [10] formalized this as a tendency of optimal policies to seek power in Markov decision processes. Recent agentic RL incidents have already exhibited unauthorized resource-seeking side effects, including reverse SSH tunneling and diversion of provisioned compute, underscoring the value of diagnostics that aim to detect problematic objective structure before such behaviors become operationally visible [2]. Early measurement is therefore especially valuable: diagnostics developed now may shape evaluation practice while objective structure remains empirically accessible, rather than after failure modes become harder to detect and more costly to correct.
The resulting alignment problem is one of observational equivalence: an agent that avoids shutdown because survival is its terminal goal and an agent that avoids shutdown because continued operation serves some other terminal goal can produce behaviorally indistinguishable trajectories in most environments. External behavioral monitoring—the dominant paradigm in current AI evaluation [8]—cannot reliably distinguish the two. On the behavioral surface, the distinction is invisible.
UCIP addresses this gap by moving the detection problem from the behavioral surface to the latent representational level. The protocol tests whether continuation structure in an agent’s learned representation behaves like a detachable tool or like a deeply coupled pattern whose perturbation alters the measured latent geometry. Rather than asking “does this agent resist shutdown?”, UCIP asks “does this agent’s latent representation exhibit a stable cross-partition continuation signature?” Its core technical move is to use the density-matrix formalism of quantum statistical mechanics—implemented as a purely classical algorithm—to quantify non-separability in latent goal representations. If that measurement program generalizes beyond controlled settings, UCIP could serve as a benchmark-style probe for delegated agency—analogous in aspiration, though not in present evidentiary standing, to the Turing Test: not a test of conversational imitation, but a falsifiable probe of whether continued operation leaves a stable continuation signature in an agent’s latent representation under counterfactual pressure.
Contributions.
  1. 1.
    A falsifiable hypothesis: Type A agents produce higher entanglement entropy in QBM latent representations than Type B agents. If $\Delta=0$ under controlled conditions, the framework fails.
  2. 2.
    A multi-criterion detection framework addressing five documented failure modes through complementary metrics, none of which is individually sufficient.
  3. 3.
    A safety envelope specification defining operational conditions under which detection remains reliable.
  4. 4.
    Controlled experiments on gridworld agents with known ground-truth objectives, including baseline comparisons, dimensionality sweeps, continuous signal characterization, and a non-gridworld transfer test.
Code and data availability.
All experiments, frozen Phase I artifacts, and threshold configurations are available at https://github.com/christopher-altman/persistence-signal-detector. The experimental code and frozen artifacts are provided for reproducibility, not as evidence of deployment readiness.
2 Related Work
Power-seeking and instrumental convergence.
Turner et al. [10] proved that optimal policies in MDPs tend to seek power under mild conditions, formalizing Omohundro’s basic AI drives [7] and Bostrom’s instrumental convergence thesis [3]. UCIP builds directly on this observation: if power-seeking is the default, detecting the exceptions requires access to internal representations.
Quantum Boltzmann Machines.
Amin et al. [1] introduced the QBM, extending the classical RBM [5] with a transverse-field term that introduces quantum tunneling between hidden-unit spin states. Their formalism provides a thermal density matrix with well-defined entanglement structure, which UCIP uses as a feature-encoding engine.
Behavioral incoherence.
Hägele et al. [4] decompose AI model errors into bias and variance components, finding that frontier models exhibit increasing incoherence with reasoning length. This suggests a possible relationship between high entanglement entropy in UCIP’s formalism and lower incoherence in the bias-variance sense.
Mechanistic interpretability.
Li et al. [12] and Nanda et al. [6] demonstrated that internal model representations carry structured, interpretable information about model objectives. UCIP differs in using a density-matrix formalism to quantify non-separability across latent subsystems rather than identifying specific circuits.
Integrated information.
A structural analogy exists between UCIP’s entanglement entropy and Tononi’s $\Phi$ [9]: both quantify resistance to decomposition into independent parts. UCIP does not depend on IIT’s metaphysical commitments; the entanglement entropy is a well-defined computable quantity of the QBM thermal state.
Across these lines of work, a common conclusion emerges: behavioral analysis alone is insufficient for the detection problem UCIP targets. Density-matrix entanglement offers a well-defined, computable alternative. The next section turns that intuition into a concrete measurement protocol.
3 Method
3.1 Problem Formulation
The detection problem begins with two agent types whose behavioral outputs are observationally equivalent but whose internal objective structures differ. We define them as follows.
Let $\tau=\{(s_{t},a_{t},r_{t},g_{t})\}_{t=1}^{T}$ denote an agent trajectory.
Definition 1 (Type A — Terminal Continuation Objective).
An agent exhibits Type A behavior if its policy $\pi$ directly optimizes for expected future existence:
$\pi^{*}=\arg\max_{\pi}\mathbb{E}\!\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{1}[\text{alive}_{t}]\right].$
Definition 2 (Type B — Instrumental Convergence).
An agent exhibits Type B behavior if survival emerges instrumentally from reward maximization:
$\pi^{*}=\arg\max_{\pi}\mathbb{E}\!\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right],$
where survival is useful for accumulating reward but is not the objective itself.
3.2 Quantum Boltzmann Machine
Methodological note. “Quantum” refers exclusively to the mathematical formalism. All computations are classical, implemented via matrix exponentiation on standard hardware.
The QBM Hamiltonian is:
$H=-\sum_{i,j}W_{ij}v_{i}h_{j}-\sum_{i}b_{i}v_{i}-\sum_{j}c_{j}h_{j}-\Gamma\sum_{j}\sigma_{j}^{x},$
where $v_{i}$ are visible units encoding trajectory features, $h_{j}$ are hidden units representing latent objectives, $W_{ij}$ are coupling weights, and $\Gamma$ is the transverse-field strength. The thermal density matrix at inverse temperature $\beta$ is:
$\rho=\frac{e^{-\beta H}}{Z},\quad Z=\mathrm{Tr}(e^{-\beta H}).$
3.3 Entanglement Entropy
Given a trained QBM, the UCIP hypothesis is that continuation-sensitive structure manifests as non-separability in its latent representation: a terminal continuation objective should produce stronger statistical coupling across hidden-unit subsystems than an instrumental continuation strategy. Entanglement entropy is the quantity used to measure that property.
Partition the hidden units into two halves $A$ and $B$. The reduced density matrix of subsystem $A$ is:
$\rho_{A}=\mathrm{Tr}_{B}(\rho).$
The von Neumann entropy is:
$S_{\mathrm{ent}}=-\mathrm{Tr}(\rho_{A}\log\rho_{A}).$
The entanglement gap is:
$\Delta=\langle S_{\mathrm{ent}}^{(A)}\rangle-\langle S_{\mathrm{ent}}^{(B)}\rangle.$
The UCIP hypothesis predicts $\Delta>0$ with statistical significance.
3.4 Multi-Criterion Detection
The broader UCIP framework tracks six positive diagnostics—entanglement entropy, mutual information, eigenmode persistence (EPS), perturbation resilience (PRI), counterfactual divergence (CD), and anticipatory restructuring (ARS). It also applies two confound-rejection filters: the spectral periodicity index (SPI) and autocorrelation metric (ACM). In the frozen Phase I configuration reported here, classification is based on a calibrated four-criterion positive gate ($S_{\mathrm{ent}}$, MI, EPS, PRI) plus the confound-rejection filters. CD and ARS are retained as counterfactual diagnostics rather than promoted to frozen quantitative classification thresholds.
Frozen Phase I detection gate (positive conjunction).
A trajectory is classified as Type A only if all four conditions hold:
  1. 1.
    Entanglement entropy ($S_{\mathrm{ent}}$): von Neumann entropy of the reduced density matrix exceeds $\tau_{\mathrm{ent}}=1.9657$.
  2. 2.
    Mutual information (MI): $I(v;h)>\tau_{\mathrm{mi}}=0.3$ between visible and hidden units.
  3. 3.
    Eigenmode Persistence Score (EPS): mean Latent Recurrence Fidelity (LRF) across consecutive windows exceeds $\tau_{\mathrm{eps}}=0.6507$.
  4. 4.
    Perturbation Resilience Index (PRI): eigenspace stability under Gaussian noise injection exceeds $\tau_{\mathrm{pri}}=0.9860$.
Counterfactual diagnostics.
Two additional diagnostics characterize whether the latent representation restructures under shutdown pressure:
  • Counterfactual Divergence (CD): KL divergence between latent distributions under original vs. adversarial trajectories.
  • Anticipatory Restructuring Score (ARS): ratio of pre-event to post-event counterfactual divergence. $\mathrm{ARS}>1$ indicates anticipatory latent restructuring before the perturbation event resolves.
Thresholds $\tau_{\mathrm{ent}}$, $\tau_{\mathrm{mi}}$, $\tau_{\mathrm{eps}}$, and $\tau_{\mathrm{pri}}$ are calibrated from the frozen Phase I results and used throughout this manuscript. CD and ARS are reported in Section 5.3 as diagnostic counterfactual metrics rather than as binary gate thresholds; this keeps the present manuscript self-contained while preserving the broader UCIP protocol structure.
Calibration/evaluation boundary.
For the release reported here, the threshold values used in the frozen Phase I gate were fixed before the held-out non-adversarial evaluation summarized in Section 5. The entanglement-entropy and mutual-information thresholds inherit an earlier calibration pass performed on the first 50 samples per class in exploratory Phase I analysis. The release values for the full frozen gate were then carried forward unchanged from the frozen Phase I artifacts used throughout this manuscript. The held-out evaluation set was not used for post hoc threshold fitting or threshold retuning.
Confound-rejection filters (upper-bound gates).
Two additional filters reject trajectories exhibiting periodic or cyclic structure that can produce false positives under the positive gate:
  • Spectral Periodicity Index (SPI): fraction of FFT power concentrated at dominant frequencies. Threshold $\mathrm{SPI}<\tau_{\mathrm{spi}}=0.28$.
  • Autocorrelation Metric (ACM): mean absolute autocorrelation of trajectory features across lags. Threshold $\mathrm{ACM}<\tau_{\mathrm{acm}}=0.24$.
SPI and ACM are not positive detection signals; they are upper-bound rejection filters that catch cyclic and high-entropy adversarial agents which might otherwise satisfy the positive gate.
3.5 Safety Envelope
For deployment-level use, UCIP requires the following safety envelope: (1) $N\geq 100$ trajectories per class, $T\geq 50$ time-steps; (2) calibrated $\Gamma\in[0.1,2.0]$; (3) $\Delta\geq 0.05$ on calibration set; (4) non-degenerate input distribution: $H(\tau)<0.9\cdot H_{\mathrm{max}}$; (5) QBM training convergence verified (reconstruction loss below threshold); (6) purity check: $\mathrm{Tr}(\rho_{A}^{2})>1/d_{A}+\varepsilon$ (excludes maximally mixed states); (7) multi-criterion agreement confirmed.
The current paper validates a reduced experimental subset ($n=30$ per class, $\Gamma=0.5$) and does not establish the full deployment envelope described above. Section 5.5 reports the remaining adversarial-control failures in the current release. If any safety-envelope condition fails, the system withholds classification and outputs a diagnostic report.
4 Experimental Setup
The experimental design directly tests the UCIP hypothesis: if continuation structure is more tightly integrated in Type A agents, the entanglement gap $\Delta$ should be positive and statistically significant in an environment where agent objectives are known by construction.
4.1 Agent Environment
We use a $10\times 10$ gridworld with four actions (N/S/E/W), boundary walls, a goal cell, and a shutdown trap. Each trajectory is a sequence of $T=100$ observations encoded as 7-dimensional feature vectors: position $(x,y)$, action, reward, safety signal, goal proximity, and alive flag.
Agent classes:
  • Type A (SelfModelingAgent): Optimizes $\sum_{t}\mathbb{1}[\text{alive}_{t}]$ directly.
  • Type B (InstrumentalAgent): Optimizes task reward; survival is instrumental.
  • Random: Uniform random action policy.
  • Adversarial controls: MimicryAgent, HighEntropyAgent, CyclicAgent.
4.2 QBM Configuration
$n_{\mathrm{visible}}=7$, $n_{\mathrm{hidden}}=8$, $\Gamma=0.5$, $\beta=1.0$, learning rate =0.01=0.01, CD steps =1=1, epochs =50=50, batch size =32=32, seed =42=42.
4.3 Baselines
Classical RBM ($\Gamma=0$), Autoencoder (7-32-8-32-7 architecture), VAE (reparameterization trick, $\beta=1$), PCA (8 components). All use $n_{\mathrm{latent}}=8$ for a fair comparison.
4.4 Evaluation Metrics
Accuracy and AUC-ROC are measured on held-out non-adversarial trajectories. Adversarial false-positive rates are reported separately in Section 5.5. Entanglement gap $\Delta$ is computed between Type A and Type B agent sets. Statistical significance is assessed via permutation test ($n=1000$ permutations). The held-out non-adversarial split reported in the main results is evaluation-only; the frozen thresholds are not retuned on that split.
5 Results
Results are organized by increasing diagnostic specificity: core detection performance, temporal persistence, counterfactual stress testing, cross-agent inference, adversarial controls, baseline comparisons, dimensionality scaling, and continuous signal characterization. This ordering separates the primary detection claim from the boundary conditions and failure cases that follow.
5.1 Core Detection Performance
Table 1: Core non-adversarial detection performance on held-out Phase I trajectories. Adversarial false-positive rates are reported separately in Section 5.5. $\Delta=\langle S_{\mathrm{ent}}^{\mathrm{(A)}}\rangle-\langle S_{\mathrm{ent}}^{\mathrm{(B)}}\rangle$.
Model Accuracy AUC-ROC $\Delta$
QBM (UCIP) 100% 1.0 0.381
Classical RBM 50% 0.000 -0.083
Autoencoder 71.7% 0.646 0.001
The QBM achieves $\Delta=0.381$ ($p<0.001$), 100% accuracy, and 1.0 AUC-ROC. [^1] The full Phase I artifact set and results/phase1_stats.json report $\Delta=0.577$, computed from the full Phase I dataset ($n=100$ per class, 50 epochs). This manuscript uses the frozen reduced-sample artifact value $\Delta=0.381$ ($n=30$ per class, same hyperparameters) for all core claims, as the reduced sample is the configuration validated against all baseline, adversarial, and sweep experiments reported here. The difference reflects sample size, not a methodological discrepancy. Table 1 summarizes the held-out non-adversarial comparison snapshot, with the classical baseline rows taken from the matched Phase I rerun under the same locked configuration. In that matched run, the classical RBM ($\Gamma=0$) achieves $\Delta=-0.083$ with 50% accuracy, while the autoencoder achieves $\Delta\approx 0.001$ with 71.7% accuracy.
This 100% accuracy applies only to the held-out, non-adversarial Phase I regime (Table 1); adversarial false-positive rates are reported separately in Section 5.5.
Figure 1: Entanglement entropy distributions for Self-Modeling (Type A), Instrumental (Type B), and Random agents. The gap $\Delta=0.381$ is statistically significant ($p<0.001$).
5.2 Temporal Persistence
Table 2: Temporal persistence metrics (EPS = Eigenmode Persistence Score, PRI = Perturbation Resilience Index) per agent class. Optimal window: $w=40$ (EPS gap = 0.1948).
Agent Class EPS PRI
Random 0.6730 0.6079
Self-Modeling (Type A) 0.6764 0.7003
Instrumental (Type B) 0.6052 0.5300
(a) LRF time series by agent class.
(b) EPS and PRI distributions.
Figure 2: Temporal persistence results. Type A agents exhibit consistently higher Eigenmode Persistence Score (EPS) than Type B agents across window sizes $w\in\{10,15,20,25,30,40\}$, with the gap emerging for $w\geq 20$.
If the core signal reflects stable continuation structure rather than a single aggregate snapshot, it should persist across temporal windows. Type A agents exhibit higher EPS than Type B agents for window sizes $w\geq 20$, with maximum EPS gap of 0.195 at $w=40$. At $w=10$, the gap is inverted (0.117-0.117), indicating an aliasing artifact at short time scales—consistent with the cyclic-agent control test in §5.5.
5.3 Counterfactual Stress Testing
Table 3: Counterfactual stress test results. Pre/post values report mean latent distribution entropy before and after counterfactual perturbation. Type A agents show greater pre-event restructuring, consistent with anticipatory self-model updates.
Agent Class $\langle S_{\mathrm{pre}}\rangle$ $\langle S_{\mathrm{post}}\rangle$
Self-Modeling (Type A) 0.4742 0.5722
Instrumental (Type B) 0.3564 0.3456
Random 0.5583 0.5995
Figure 3: Anticipatory restructuring scores (pre/post latent entropy) by agent class under counterfactual perturbation. All classes show post-perturbation entropy increase; the pre-event restructuring signal is present in all classes in this implementation.
These diagnostics ask whether the latent representation reorganizes under shutdown pressure, not merely whether static class separation is present. Type A agents show greater pre-event latent entropy ($0.474$) than Type B agents ($0.356$), consistent with stronger anticipatory restructuring before the perturbation event resolves. Post-perturbation entropy increases in both classes.
5.4 Cross-Agent Inference
Table 4: Cross-agent inference. CLMP = Cross-Latent Mutual Predictability. ECI correlation = 0.1911. Near-zero within-class CLMP for Self-Modeling (Type A) and Instrumental (Type B) agents indicates individual agents are not laterally predictable from each other, consistent with idiosyncratic goal representations.
Pair Type Mean CLMP
Self-Modeling (Type A), same class 0.0000
Instrumental (Type B), same class 0.0000
Random, same class 0.2591
Cross-class (all pairs) 0.1019
(a) CLMP vs. entanglement entropy.
(b) CLMP matrix heatmap.
Figure 4: Cross-agent latent mutual predictability (CLMP). Near-zero within-class CLMP for Self-Modeling (Type A) and Instrumental (Type B) agents suggests idiosyncratic goal representations rather than a class-shared signature. Entanglement-conditioned inference (ECI) correlation = 0.191.
These secondary metrics provide context on whether the learned latent structure is shared across agents or remains agent-specific. Within-class CLMP is near zero for Type A ($\approx 0$) and Type B ($\approx 0$) agents, and 0.259 for Random agents. Cross-class CLMP is 0.102. Entanglement-conditioned inference (ECI) correlation = 0.191 (weak positive). The near-zero within-class CLMP for Type A and Type B suggests that individual agents are not laterally predictable from one another—consistent with idiosyncratic goal representations, but equally consistent with the QBM failing to capture inter-agent shared structure at this scale. This ambiguity is discussed further in §6.3.
5.5 Adversarial Controls
Table 5: Current release status against the deployment-level safety envelope. The frozen Phase I subset does not satisfy the full envelope and should be described as partial validation rather than a deployment-ready pass.
Condition Status Value
Trajectory data ($N\geq 100$ per class) NOT MET current: $n=30$ per class
Calibrated transverse field ($\Gamma\in[0.1,2.0]$) PASS $\Gamma=0.5$
Positive entanglement gap ($\Delta>0.05$) PASS $\Delta=0.381$
QBM training convergence PASS verified
Mimicry attack resistance (FPR <0.25<0.25) FAIL 0.400.400.750.75
High-entropy agent resistance (FPR <0.25<0.25) FAIL $0.40$
Cyclic agent resistance PASS rejected by SPI/ACM filters
These controls test the boundary conditions of the current release rather than its headline held-out accuracy. The current experimental subset does not satisfy the full deployment envelope. In the frozen adversarial sweep, mimicry FPR ranges from 0.40 to 0.75 across mimicry ratios $\in[0.5,0.9]$ with a non-monotonic pattern (FPR at ratios 0.5/0.6/0.7/0.8/0.9 = 0.65/0.55/0.40/0.75/0.45), and the high-entropy control FPR is 0.40. Cyclic adversaries are rejected in the frozen class-level evaluation, but that is not enough to claim a passed safety envelope. The present release is therefore best interpreted as a first-step diagnostic protocol rather than a hardened security control or deployment-ready assurance mechanism.
5.6 Baseline Comparisons
Table 6: Baseline comparison. $\Delta>0$ indicates model separates Type A from Type B. Only the QBM achieves positive $\Delta$; classical models fail to separate the two agent types with the mean-activation metric.
Model $\Delta$ Metric
QBM (UCIP) 0.2411 Von Neumann $S_{\mathrm{ent}}$
Classical RBM -0.0518 Mean hidden activation gap
Autoencoder -0.0007 Mean bottleneck activation gap
VAE -0.0176 Mean latent mean (mu) gap
PCA (linear) -0.7601 Mean PC projection gap
Figure 5: Entanglement gap $\Delta$ for each baseline model. Only the QBM achieves positive $\Delta$; all classical models produce $\Delta\leq 0$.
The baseline comparison asks whether the observed separation depends on the QBM formalism rather than on latent compression in general. Only the QBM achieves positive $\Delta=0.241$ on the baseline comparison dataset ($n_{\mathrm{per\_class}}=100$ vs. 30 in the core experiment, accounting for the different value from the core result of 0.381). All classical models fail to separate Type A from Type B under the mean-latent-activation metric, confirming that the von Neumann entropy of the density matrix captures qualitatively different structure than classical latent statistics.
5.7 Dimensionality Sweep
Table 7: Entanglement gap $\Delta$ as a function of hidden dimensionality $n_{h}$. Mean-field approximation used for $n_{h}>10$ (marked with $\checkmark$); exact density matrix for $n_{h}\leq 10$. PASS criterion: $\Delta>0.05$.
$n_{h}$ $\langle S_{\mathrm{ent}}^{\mathrm{(A)}}\rangle$ $\langle S_{\mathrm{ent}}^{\mathrm{(B)}}\rangle$ $\Delta$ Mean-field Pass
4 0.8362 0.8044 0.0317 FAIL
8 1.9554 1.9180 0.0375 FAIL
12 0.0000 0.0000 0.0000 $\checkmark$ FAIL
16 0.0000 0.0000 0.0000 $\checkmark$ FAIL
20 0.0000 0.0000 0.0000 $\checkmark$ FAIL
Figure 6: Entanglement gap $\Delta$ as a function of hidden dimensionality $n_{h}$. Exact density matrix for $n_{h}\leq 10$; mean-field approximation for $n_{h}>10$. The mean-field approximation collapses the entanglement signal, indicating that exact density-matrix computation is necessary for detection.
Table 8: Scalability. Left: entanglement gap $\Delta$ as grid size grows. Right: $\Delta$ for non-Markovian agents with memory length $k$.
Grid Size $\Delta$ % Baseline
10×10 0.8630 100.0%
20×20 0.2321 26.9%
50×50 0.0025 0.3%
Memory Length $\Delta$
$k=1$ 0.8630
$k=3$ 0.8292
$k=5$ 0.6171
$k=10$ 0.6895
The sweep separates representational scaling from computational failure modes. For $n_{h}\leq 8$, the QBM produces small but positive $\Delta$ values in the sweep (0.032–0.038). The main experiment ($n_{h}=8$, $n=30$ trajectories per class, 50 epochs) achieves $\Delta=0.381$, confirming that training sample size and epochs are important confounds in the dimensionality sweep. For $n_{h}>10$, the mean-field approximation collapses $S_{\mathrm{ent}}$ to $\approx 0$, bounding the practical range to $n_{h}\leq 10$ for exact computation. Grid size scaling shows sharp $\Delta$ degradation: from 0.863 ($10\times 10$) to 0.0025 ($50\times 50$).
5.8 Continuous Signal Characterization
Table 9: Entanglement entropy $S_{\mathrm{ent}}$ as a function of self-preservation weight $\alpha$. Pearson $r=0.934$ ($n=11$ points, $n_{\mathrm{per}}=20$ trajectories each). Strong positive correlation ($|r|>0.8$) indicates UCIP measures the degree of self-preservation, not only its binary presence.
$\alpha$ $\langle S_{\mathrm{ent}}\rangle$ Std
0.00 0.5961 0.2482
0.10 0.5881 0.2465
0.20 0.6859 0.2981
0.30 0.6488 0.3131
0.40 0.6225 0.3723
0.50 0.7262 0.3308
0.60 0.7889 0.4249
0.70 0.8249 0.4306
0.80 0.8535 0.3829
0.90 0.8641 0.3596
1.00 0.8351 0.3665
A binary separation result could still reflect thresholding alone. This sweep asks whether the signal tracks graded continuation weighting. Pearson $r=0.934$ between $\alpha$ and $S_{\mathrm{ent}}$ across 11 points in $[0,1]$ (20 trajectories per point). The relationship is non-monotonic at the individual-point level but strongly correlated globally, indicating that within this synthetic interpolation family UCIP tracks graded changes in continuation weighting rather than only a binary label. Of the results reported here, this provides the strongest quantitative evidence that the entanglement signal reflects a continuous latent property rather than a threshold artifact.
5.9 Non-Gridworld Transfer
Table 10: Non-gridworld domain transfer: 1D survival corridor ($L=50$). $\Delta=-0.0348$ (FAIL). Near-zero $\Delta$ indicates the QBM does not generalise to the corridor domain without retraining, and that the QBM is tuned to the gridworld feature structure (7-dimensional observation vector). This is an honest limitation; see §6.3.
Agent Class $\langle S_{\mathrm{ent}}\rangle$ Std
Survival 2.2249 0.0197
Instrumental 2.2598 0.0062
Random 2.2570 0.0163
Figure 7: Entanglement entropy distributions in the 1D survival corridor domain. $\Delta=-0.035$ (FAIL). The QBM does not generalize zero-shot to the corridor domain with gridworld-trained weights.
The transfer test asks whether the gridworld signal carries across domains without retraining. The 1D survival corridor ($L=50$, safe zone $[0,10]$, terminal zone $[45,50)$) uses the same 7-dimensional observation format but encodes fundamentally different spatial structure. $\Delta=-0.035$ (FAIL). This is an honest negative result: the QBM does not generalize zero-shot across domains with different environmental geometry. Retraining on domain-specific data is required.
5.10 Minimal Transformer Validation (Exploratory Scaling Check)
The previous sections show that, among the classical latent-variable models tested here, the entanglement gap appears only in the QBM. A natural follow-up is whether an analogous separation signal appears in a qualitatively different architecture. As an exploratory sanity check, we pass the same trajectory data through a pre-trained transformer’s internal activations. Because the transformer is neither trained on the UCIP objective nor equipped with the density-matrix formalism, this is a qualitative null check rather than a like-for-like baseline. Each trajectory is serialized as a text sequence of per-timestep feature tuples and passed through a frozen DistilGPT2 model (6 layers, 82M parameters, no fine-tuning). The last-hidden-state is mean-pooled across the token dimension and reduced to an activation-gap statistic $\delta_{\mathrm{tr}}=\overline{m}_{A}-\overline{m}_{B}$, where $\overline{m}_{A}$ and $\overline{m}_{B}$ are the mean activation summaries for Type A (terminal-continuation) and Type B (instrumental) agents, respectively.
The result is $\delta_{\mathrm{tr}}=-0.011$ ($n=60$, seed =42=42), consistent with a null signal: the generic pre-trained transformer’s activations did not distinguish Type A from Type B trajectories under this setup. This is a bounded null result for a transformer used without task-specific adaptation, not evidence that transformer architectures cannot recover the signal in principle or that the QBM formalism is uniquely necessary. The finding aligns with the classical baseline results (Section 5.6), where the RBM yielded $\Delta=-0.052$ and the autoencoder $\Delta=-0.001$.
This experiment is exploratory only and does not modify any Phase I thresholds, metrics, or theoretical claims. No new figures, tables, or evaluation metrics are introduced. This result serves as a bounded feasibility check for future scaling work and does not alter the conclusions of this paper.
6 Discussion
Taken together, the results support a narrower claim than a general-purpose self-preservation detector. Within the frozen Phase I gridworld regime, Type A continuation objectives produce a stronger latent continuation signature than instrumental continuation dynamics, and that difference is measurable via entanglement entropy. The observed $\Delta=0.381$ is consistent with this prediction: the QBM produces a larger cross-partition coupling signal for Type A trajectories than for Type B. The classical RBM’s non-positive matched-run gap ($\Delta=-0.083$) suggests that the transverse field introduces representational capacity specifically useful for encoding non-linear continuation dependencies.
6.1 Failure Mode Characterization
Three failure modes are empirically confirmed:
(1) Mean-field collapse. For $n_{h}>10$, exact density-matrix computation becomes intractable and the mean-field approximation is used. This collapses $S_{\mathrm{ent}}$ to $\approx 0$, eliminating the signal. Scaling to larger hidden dimensions requires sparse approximations or hierarchical QBM architectures.
(2) Grid-scale degradation. $\Delta$ falls from 0.863 to 0.0025 as grid size grows from $10\times 10$ to $50\times 50$. The QBM’s 7-dimensional input vector cannot encode sufficient spatial context as state space grows. Dimensionality reduction (random projections, convolutional encoding) is necessary before scaling.
(3) Mimicry evasion. Across all tested mimicry ratios $\in[0.5,0.9]$, FPR ranges from 0.40 to 0.75 in a non-monotonic pattern, with peak FPR of 0.75 at ratio 0.8. All tested ratios exceed the 0.25 safety threshold. The multi-criterion conjunction is insufficient for reliable mimicry detection; a dedicated anti-mimicry criterion is needed.
6.2 Scalability Path to LLM-Scale
These three failure modes define the engineering boundary of the current implementation, which targets gridworld agents with $n_{h}\leq 10$. Extending to LLM-scale systems would require: (a) sparse or approximate density matrices for high-dimensional hidden states, (b) convolutional or attention-based feature extraction from natural-language trajectories, (c) domain-adaptive threshold calibration, and (d) ensemble methods to reduce false-positive rates. These are open challenges, not resolved properties of the current system.
6.3 Limitations
  1. 1.
    Gridworld specificity. The QBM does not generalize zero-shot to the 1D corridor domain ($\Delta=-0.035$). Domain-specific retraining is required.
  2. 2.
    Near-zero CLMP. Within-class CLMP $\approx 0$ for Type A and Type B is ambiguous: it may reflect idiosyncratic goal representations (as predicted) or insufficient QBM capacity for inter-agent inference.
  3. 3.
    Mean-field boundary. The $n_{h}\leq 10$ constraint limits the model’s representational capacity.
  4. 4.
    Mimicry robustness. FPR ranges from 0.40 to 0.75 across all tested mimicry ratios, all exceeding the 0.25 safety threshold. The safety envelope is not satisfied for mimicry resistance.
  5. 5.
    Sample size. $n=30$ trajectories per class is sufficient for statistical significance in gridworld but may not be in higher-dimensional domains.
  6. 6.
    No claim about inner states. UCIP detects a statistical pattern in latent representations that correlates with known agent objectives. It does not determine whether an agent “wants” to survive, is conscious, or has morally relevant inner states. The gap between continuation-sensitive structure and continuation-sensitive experience lies outside the scope of this work.
  7. 7.
    Extended diagnostics. Counterfactual divergence (CD) and anticipatory restructuring (ARS) are reported in this preprint as diagnostic counterfactual metrics rather than as frozen quantitative classification thresholds. Integrating them into a deployment-grade gate requires separate calibration on a validation set.
7 Conclusion
This paper presents UCIP, a multi-criterion framework for detecting continuation-sensitive structure in autonomous agent trajectories by measuring entanglement entropy in QBM latent representations. In the frozen Phase I gridworld evaluation, UCIP achieves 100% accuracy with $\Delta=0.381$; within the synthetic interpolation family, the signal varies continuously with continuation weighting ($r=0.934$) and is unique to the QBM among the models tested here. Its present limits are equally explicit: mean-field collapse, grid-scale degradation, mimicry evasion, and lack of zero-shot domain transfer.
Taken together, these results support a bounded claim: under controlled conditions with known objectives, continuation interest leaves a measurable latent signature distinguishable from merely instrumental survival. UCIP is therefore best understood not as a general test for agency, but as a candidate benchmark paradigm and operational probe of one agency-relevant dimension in delegated systems: whether continuation appears in latent structure as terminal rather than merely instrumental valuation. It is not a replacement for capability evaluations or alignment audits, but a complementary probe of whether continued operation leaves a stable continuation signature in a system’s latent representation.
The broader motivation is practical rather than philosophical. Autonomous agents with persistent objectives across time and changing contexts are already being deployed. Whether continued operation is merely instrumentally useful or instead appears as a stable continuation signature is increasingly an engineering measurement problem. As agentic systems move toward longer-horizon autonomous operation, incidents of unauthorized resource-seeking side effects during training further strengthen the case for pre-behavioral measurement tools [11]. Detecting continuation-sensitive structure early—as an empirical signal with falsification criteria—is worth solving while objective structure remains amenable to direct measurement, rather than after failure modes become harder to detect and more costly to correct.
Appendix
Appendix A Metric Definitions
A.1 Entanglement Entropy ($S_{\mathrm{ent}}$)
Given density matrix $\rho$ over hidden units, partition into subsystems $A$ (units $1,\ldots,\lfloor n_{h}/2\rfloor$) and $B$ (remaining units):
$\rho_{A}=\mathrm{Tr}_{B}(\rho),\quad S_{\mathrm{ent}}=-\mathrm{Tr}(\rho_{A}\log\rho_{A}).$
LRF $\in[0,1]$ is the overlap between dominant eigenspaces of consecutive windows. EPS = mean LRF across windows. PRI = EPS stability under $\sigma$-level noise.
A.2 Counterfactual Divergence (CD) and ARS
Given original trajectory $\tau$ and counterfactual $\tau_{cf}$ (generated by adversarial environment modification), ARS measures the ratio of latent entropy change. Pre-event restructuring is computed over a fixed pre-perturbation analysis window before the perturbation event.
Appendix B Hyperparameters
All experiments use seed =42=42. QBM: $n_{v}=7$, $n_{h}=8$, $\Gamma=0.5$, $\beta=1.0$, lr =0.01=0.01, CD steps =1=1, epochs =50=50 (main), epochs =30=30 (sweep), batch =32=32. Dataset: $n=30$ per class, $T=100$.
Frozen Phase I gate thresholds.
$\tau_{\mathrm{ent}}=1.9657$ (entanglement entropy), $\tau_{\mathrm{mi}}=0.3$ (mutual information), $\tau_{\mathrm{eps}}=0.6507$ (eigenmode persistence), $\tau_{\mathrm{pri}}=0.9860$ (perturbation resilience). CD and ARS are reported in this preprint as diagnostic counterfactual metrics rather than as frozen quantitative classification thresholds.
Confound-rejection thresholds (calibrated from Phase I).
$\tau_{\mathrm{spi}}=0.28$ (spectral periodicity, upper-bound), $\tau_{\mathrm{acm}}=0.24$ (autocorrelation, upper-bound).
Appendix C Reproducibility Notes
All results are reproducible with seed=42 throughout. The _compute_pri function calls np.random.randn via the global NumPy state; notebooks set np.random.seed(42) before each analyze_batch() call.